OpenStudy (x3_drummerchick):

how do i find the asymptotes of the hyperbola shown below? Will give medals!

OpenStudy (x3_drummerchick):

\[\frac{ (x-1)^{2} }{ 5 } - \frac{ (y-2)^{2} }{ \frac{ 5 }{ 4 } }\]

OpenStudy (x3_drummerchick):

i know that the asymptotes= plus/minus b/a, but things got weird due to the shift

OpenStudy (x3_drummerchick):

\[y=\frac{ 5 }{ 2 }-\frac{ x }{ 2 } and y=\frac{ 3 }{ 2 }+\frac{ x }{ 2 }\]

OpenStudy (x3_drummerchick):

those are the answers, but i dont understand how they are gotten?

OpenStudy (x3_drummerchick):

@dan815

OpenStudy (x3_drummerchick):

@mathmale

OpenStudy (x3_drummerchick):

@zepdrix

OpenStudy (x3_drummerchick):

@xapproachesinfinity